OpenStudy (anonymous):

If x is a number that is 7 more than y, how do you express x as a function of y?

OpenStudy (anonymous):

what about if x was 3 times more than y i would say 3x = y

OpenStudy (anonymous):

if x is a number 22 less than y i would say x - 22 = y

OpenStudy (anonymous):

my answer choice are only equations

OpenStudy (anonymous):

x=f(y)=7+x x=f(y)=7+y y=f(x)=7x y=f(x)=y-7 y=f(x)=7+y

OpenStudy (anonymous):

what do you think it is. i have gave you examples of how to work them in the above text

OpenStudy (anonymous):

i dont know. i didnt really much understand what you meant

Nnesha (nnesha):

divide that into separate part change into math more than meaning what sign you should use ?? \(\large\color{red}{is}\) meaning "=" equal sign

OpenStudy (dan815):